Find the coordinates of the vertices, even if they are not part of the solution
y-3<0
y<2x+1
3y-x>=3

Answer by stanbon(75887)

Find the coordinates of the vertices, even if they are not part of the solution
y-3<0
y<2x+1
3y-x>=3
----
Solve each for "y"::
y < 3
y < 2x+1
y > x/3 + 1
-----------------
Solve each pair of EQUALITIES to find the points of intersection::
3 = 2x+1
x = 1, then y = 3------- vertex (1,3)
-----
3 = x/3 + 1
x = 6, then y = 3 ------ vertex (6,3)
-----
2x+1 = x/3+1
(5/3)x = 0
x = 0, then y = 1 ------ vertex (0,1)
------------------------
Cheers,
Stan H.